# Viscoelastic Worthington jets & droplets produced by bursting bubbles

### v1

@@ Abstract
Bubble bursting and subsequent collapse of the open cavity at free surfaces of contaminated liquids can generate aerosol droplets, facilitating pathogen transport. After film rupture, capillary waves focus at the cavity base, potentially generating fast Worthington jets that are responsible for ejecting the droplets away from the source. While extensively studied for Newtonian fluids, the influence of non-Newtonian rheology on this process remains poorly understood. Here, we employ direct numerical simulations to investigate the bubble cavity collapse in viscoelastic media, such as polymeric liquids, examining how their elastic modulus $G$ and their relaxation time $\lambda$ affect jet and droplet formation. We show that the viscoelastic liquids yield Newtonian-like behavior as either parameter $G$ or $\lambda$ approaches zero, while increasing them suppresses jet formation due to elastic resistance to elongational flows. Intriguingly, for some cases with intermediate values of $G$ and $\lambda$, smaller droplets are produced compared to Newtonian fluids, potentially enhancing aerosol dispersal. By mapping the phase space spanned by the elastocapillary number (dimensionless $G$) and the Deborah number (dimensionless $\lambda$), we reveal three distinct flow regimes: (i) jets forming droplets, (ii) jets without droplet formation, and (iii) absence of jet formation. Our results elucidate the mechanisms underlying aerosol suppression versus fine spray formation in polymeric liquids, with implications for pathogen transmission and industrial processes involving viscoelastic fluids.
@@ 1 Introduction
Here, $g$ is the acceleration due to gravity. The solvent Ohnesorge number $Oh_{s}$ exemplifies the dimensionless viscosity of the surrounding medium, significantly influencing the capillary wave dynamics, determining their damping and overall viscous dissipation, while the Bond number $Bo$ affects the initial cavity shape and the hydrostatic pressure differences (Walls et al., 2015; Bergmann et al., 2006, 2009; Lohse, 2018). In this study, we will focus our attention on the limiting case of very small bubbles with $Bo=0.001$, for which the bubbles can be approximated as spheres (figures 1a, Toba, 1959; Princen, 1963; Lhuissier & Villermaux, 2012). For the Newtonian cases, appendix A summarizes the key results, including the effect of $Oh_{s}$ on bubble-busting dynamics. For the influence of gravity on the shape and consequently the overall dynamics, we refer the readers to Toba (1959); Princen (1963); Walls et al. (2015); Krishnan et al. (2017); Deike et al. (2018).
@@
Figure 1: (a-i) A bubble with radius $R_{0}$ rests close to the liquid-gas interface, separated from it by a thin liquid film of thickness $\delta\ll R_{0}$. The surrounding viscoelastic medium is characterized by density $\rho_{s}$, solvent viscosity $\eta_{s}$, elastic modulus $G$, and relaxation time $\lambda$. The gas has density $\rho_{g}$ and viscosity $\eta_{g}$. (a-ii) Film rupture creates an axisymmetric cavity, which we study in this work. (b) Apart from the solvent Ohnesorge number $Oh_{s}=\eta_{s}/\sqrt{\rho_{s}\gamma R_{0}}$ and the Bond number $Bo=\rho_{s}gR_{0}^{2}/\gamma$, the presence of polymers introduces two additional parameters, namely the elastocapillary number $Ec=GR_{0}/\gamma$ (equation (3)) and the Deborah number $De=\lambda/\sqrt{\rho_{s}R_{0}^{3}/\gamma}$ (equation (4)). To explore the dynamics, we move across the entire $Ec$-$De$ phase space. Often, the polymeric Ohnesorge number $Oh_{p}=G\lambda/\sqrt{\rho_{s}\gamma R_{0}}{{\color[rgb]{0,0,0}=Ec\times De}}$ (equation (5)) based on polymeric viscosity is also used to describe the influence of polymers.
@@
Given the potential for jet drops to transport pathogens or pollutants into the atmosphere, strategies to prevent their generation are pertinent. Recent studies unsurprisingly show that non-Newtonian effects, particularly viscoelasticity, can suppress jet drop production (Sanjay et al., 2021; Sen et al., 2021; Rodríguez-Díaz et al., 2023; Ji et al., 2023). While computational studies have successfully reproduced experimental observations, such as elasticity-induced droplet suppression (Cabalgante-Corralesa et al., 2024; Balasubramanian et al., 2024), the full impact of these effects on bubble-bursting dynamics remains elusive. Advancements in solving non-linear constitutive equations for highly deformed interfacial flows have been made possible by techniques like the log-conformation method (Fattal & Kupferman, 2004) and the square-root conformation method (Balci et al., 2011). Originally developed for single-phase flows, these methods have been extended to multiphase flows (Fraggedakis et al., 2016; López-Herrera et al., 2019; Varchanis & Tsamopoulos, 2022; França et al., 2024; Zinelis et al., 2023), facilitating more comprehensive investigations into this topic.
@@
$Oh_{p}=\frac{\eta_{p}}{\sqrt{\rho_{s}\gamma R_{0}}}={{\color[rgb]{0,0,0}Ec\times De}},$ (5)
which is the product of $Ec$ and $De$.
In this study, we investigate viscoelastic effects on bubble bursting dynamics by exploring the three-dimensional phase space of $Oh_{s}$, $Ec$, and $De$, using volume of fluid-based finite volume simulations. Our results demonstrate that the addition of polymers significantly influences the overall dynamics, which are governed by the interplay of viscous and elastic effects. For systems with a permanent memory of its initial state and subsequent deformations, i.e., when the additional polymeric stresses are sustained throughout the process time scale ($De\to\infty$), the dimensionless elastic modulus dictates the dynamics and suppression of jet and drops. In contrast, for systems with poor memory of its initial state and subsequent deformation ($De\to 0$), the dynamics resemble those encountered in Newtonian liquids with an effective viscosity deduced using the slender elastic jet equations.
Building upon the extensive literature on viscoelastic flows, we extend these concepts to the specific case of bubble bursting. Previous research has explored viscoelastic phenomena in various contexts, including flow through nozzles and contractions (Hinch, 1993; Chen, 1991; Boyko et al., 2024), stability and breakup of viscoelastic jets (Middleman, 1965; Goren & Gottlieb, 1982; Bousfield et al., 1986; Chang et al., 1999; Anna & McKinley, 2001; Pandey et al., 2021; Sen et al., 2024; Zinelis et al., 2023), coalescence and spreading of viscoelastic drops and bubbles (Bouillant et al., 2022; Dekker et al., 2022; Oratis et al., 2023), and oscillating bubbles in viscoelastic media (Oratis et al., 2024). Recent studies have also investigated elastoviscoplastic flows, incorporating viscous, elastic, and plastic aspects (Putz & Burghelea, 2009; Varchanis et al., 2019; França et al., 2024; Balasubramanian et al., 2024), further expanding our understanding of non-Newtonian liquids. We refer readers to reviews by Bogy (1979), Eggers (1997), and Yarin (1993) for comprehensive overviews of these topics. Our work applies the foundational knowledge developed in these works to elucidate how viscoelasticity alters the formation of Worthington jets and ejected droplets during bubble bursting, enhancing our understanding of this specific phenomenon.
@@ 2.1 Governing Equation
with $\boldsymbol{\mathcal{D}}=\left(\boldsymbol{\nabla u}+\left(\boldsymbol{\nabla u}\right)^{T}\right)/2$ representing the symmetric part of the velocity gradient tensor. The non-Newtonian contribution $\boldsymbol{\sigma_{p}}$ arises from the presence of polymers in the fluid. We emphasize that although we refer to $\boldsymbol{\sigma_{p}}$ as ‘polymeric stresses’ in the context of dilute polymer liquids, this concept extends to any deformable microstructure within the fluid that responds to flow (Saramito, 2007; Snoeijer et al., 2020; França et al., 2024; Balasubramanian et al., 2024). To characterize the deformation of these microstructures, we introduce the conformation tensor $\boldsymbol{\mathcal{A}}$, an order parameter that evolves from an initial identity state $\boldsymbol{\mathcal{A}}=\boldsymbol{\mathcal{I}}$ (figure 1a-ii). For the stress-strain constitutive relationship, we employ the model based on the Oldroyd-B family of models, which represents the simplest conformation tensor-based approach for viscoelastic fluids (Oldroyd, 1950; Bird et al., 1977; Snoeijer et al., 2020; Stone et al., 2023; Boyko & Stone, 2024). This model assumes a linear relationship between elastic stresses and polymeric deformation,
@@
Additionally, the conformation tensor $\boldsymbol{\mathcal{A}}$ relaxes to its base state $\boldsymbol{\mathcal{I}}$ over time due to thermal effects. Once more, using the Oldroyd-B model, $\boldsymbol{\mathcal{A}}$ follows a linear relaxation law,
@@
${{\boldsymbol{\mathcal{A}}}}\equiv\frac{\partial\boldsymbol{\mathcal{A}}}{\partial t}+\left(\boldsymbol{u\cdot\nabla}\right)\boldsymbol{\mathcal{A}}-2\text{Sym}\left(\boldsymbol{\mathcal{A}\cdot}\left(\boldsymbol{\nabla u}\right)\right)$ (11)
is the frame-invariant upper convected derivative of second-rank tensor $\boldsymbol{\mathcal{A}}$, and $De=\lambda/\tau_{\gamma}$ (defined in equation (4)) is the Deborah number, representing the ratio of the polymer relaxation time $\lambda$ to the process timescale $\tau_{\gamma}$ and Sym depicts the symmetric part of the tensor. The Deborah number characterizes the polymeric liquid’s memory. It is instructive to note that in the limit of $De\to\infty$, polymeric liquids have permanent memory and the dissolved polymers undergo affine motion (see equation (10) and Snoeijer et al., 2020; Stone et al., 2023; Boyko & Stone, 2024)
@@
The Oldroyd-B model, despite its widespread use due to its simplicity, fails to capture several important physical phenomena (Snoeijer et al., 2020). It inadequately describes shear-thinning behavior in polymeric liquids (Yamani & McKinley, 2023) and erroneously predicts unbounded stress growth in strong extensional flows (McKinley & Sridhar, 2002; Eggers et al., 2020). These limitations can be addressed by incorporating finite polymer extension, for example, by increasing the effective $Ec$ as the polymer approaches full extension (Hinch & Harlen, 2021; Zinelis et al., 2023). Various extensions of the Oldroyd-B equations have been developed to account for such nonlinearity, either in equations (9) and (10) or in the solvent contribution in equation (8) (Tanner, 2000). In this study, we employ the Oldroyd-B model to include the two primary effects of the polymer addition: the additional stress ($Ec$) and polymeric liquid memory ($De$). Our aim is to provide a comprehensive understanding of the entire $Ec$-$De$ parameter space (figure 1b).
@@ 2.2 Methods
We employ the open-source software Basilisk C (Popinet & collaborators, 2013–2024; Popinet, 2015) to solve the governing equations outlined in § 2.1. The equations are solved using the one-fluid approximation (Tryggvason et al., 2011), with surface tension incorporated as singular body force at the liquid-gas interface (Brackbill et al., 1992). To account for the gas phase, in addition to the dimensionless parameters described in § 1 and § 2.1, we maintain constant density and viscosity ratios of $\rho_{r}=\rho_{g}/\rho_{s}=10^{-3}$ and $\eta_{r}=\eta_{g}/\eta_{s}=2\times 10^{-2}$, respectively. The liquid-gas interface is tracked using the volume of fluid (VoF) method, governed by the advection equation
@@
The initial bubble shape is determined by solving the Young-Laplace equations for quasi-static equilibrium (Princen, 1963; Toba, 1959; Villermaux et al., 2022; Sanjay, 2022). While the shape’s asymmetry increases with the Bond number $Bo$, we focus on the limit $Bo\to 0$, setting $Bo=0.001$ to regularize the singularity at the sphere-plane intersection. This results in a near-spherical initial cavity shape (figure 1a-i). During the bubble cap bursting, the film cap retracts almost instantaneously, after which the capillary waves are generated. As we are interested only in the bubbe cavity collapse, the simulations begin with an open cavity without the thin cap (figure 1a-ii), as also done similarly in recent studies (Deike et al., 2018; Gordillo & Rodríguez-Rodríguez, 2019; Sanjay et al., 2022). The computational domain spans $8R_{0}\times 8R_{0}$, discretized using quadtree grids with adaptive mesh refinement (AMR) (Popinet, 2009). Error tolerances for the VoF color function, curvature, velocity, and order parameter $\boldsymbol{\mathcal{A}}$ are set to $10^{-3}$, $10^{-6}$, $10^{-3}$, and $10^{-3}$, respectively.
In this work, following our earlier study (Sanjay et al., 2022), most simulations maintain a minimum grid size of $\Delta=R_{0}/512$, which dictates that, to get consistent results, 512 cells are required across the bubble radius while using uniform grids. We have also used an increased resolution ($\Delta=R_{0}/1024$ for high $De$ cases and $\Delta=R_{0}/2048$ near transitions) as needed. We have carried out extensive grid independence studies to ensure that changing the grid size does not influence the results. We refer the readers to Popinet (2015); Sanjay (2022); Sanjay & Dixit (2024) for further details of the numerical method used in this work.

### v2

@@ Abstract
Bubble bursting and subsequent collapse of the open cavity at free surfaces of contaminated liquids can generate aerosol droplets, facilitating pathogen transport. After film rupture, capillary waves focus at the cavity base, potentially generating fast Worthington jets that are responsible for ejecting the droplets away from the source. While extensively studied for Newtonian fluids, the influence of non-Newtonian rheology on this process remains poorly understood. Here, we employ direct numerical simulations to investigate the bubble cavity collapse in viscoelastic media, such as polymeric liquids. We find that the jet and drop formation are dictated by two dimensionless parameters: the elastocapillary number $Ec$ (the ratio of the elastic modulus and the Laplace pressure) and the Deborah number $De$ (the ratio of the relaxation time and the inertio-capillary timescale). We show that for low values of $Ec$ and $De$, the viscoelastic liquid adopts a Newtonian-like behavior, where the dynamics are governed by the solvent Ohnesorge number $Oh_{s}$ (the ratio of visco-capillary and inertio-capillary timescales). In contrast, for large values $Ec$ and $De$, the enhanced elastic stresses completely suppress the formation of the jet. For some cases with intermediate values of $Ec$ and $De$ , smaller droplets are produced compared to Newtonian fluids, potentially enhancing aerosol dispersal. By mapping the phase space spanned by $Ec$, $De$, and $Oh_{s}$ , we reveal three distinct flow regimes: (i) jets forming droplets, (ii) jets without droplet formation, and (iii) absence of jet formation. Our results elucidate the mechanisms underlying aerosol suppression versus fine spray formation in polymeric liquids, with implications for pathogen transmission and industrial processes involving viscoelastic fluids.
@@ 1 Introduction
Here, $g$ is the acceleration due to gravity. The solvent Ohnesorge number $Oh_{s}$ exemplifies the dimensionless viscosity of the surrounding medium, significantly influencing the capillary wave dynamics, determining their damping and overall viscous dissipation, while the Bond number $Bo$ affects the initial cavity shape and the hydrostatic pressure differences (Walls et al., 2015; Bergmann et al., 2006, 2009; Lohse, 2018). In this study, we will focus our attention on the limiting case of very small bubbles with $Bo=0.001$, for which the bubbles can be approximated as spheres (figures 1a, Toba, 1959; Princen, 1963; Lhuissier & Villermaux, 2012). For the Newtonian cases, appendix A summarizes the key results, including the effect of $Oh_{s}$ on bubble-busting dynamics. For the influence of gravity on the shape and consequently the overall dynamics of Newtonian fluids,   we refer the readers to Toba (1959); Princen (1963); Walls et al. (2015); Krishnan et al. (2017); Deike et al. (2018).
@@
Figure 1: (a-i) A bubble with radius $R_{0}$ rests close to the liquid-gas interface, separated from it by a thin liquid film of thickness $\delta\ll R_{0}$. The surrounding viscoelastic medium is characterized by density $\rho_{s}$, solvent viscosity $\eta_{s}$, elastic modulus $G$, and relaxation time $\lambda$. The gas has density $\rho_{g}$ and viscosity $\eta_{g}$. (a-ii) Film rupture creates an axisymmetric cavity, which we study in this work. (b) Apart from the solvent Ohnesorge number $Oh_{s}=\eta_{s}/\sqrt{\rho_{s}\gamma R_{0}}$ and the Bond number $Bo=\rho_{s}gR_{0}^{2}/\gamma$, the presence of polymers introduces two additional parameters, namely the elastocapillary number $Ec=GR_{0}/\gamma$ (equation (3)) and the Deborah number $De=\lambda/\sqrt{\rho_{s}R_{0}^{3}/\gamma}$ (equation (4)). To explore the dynamics, we move across the entire $Ec$-$De$ phase space. Often, the polymeric Ohnesorge number $Oh_{p}=G\lambda/\sqrt{\rho_{s}\gamma R_{0}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=Ec\times De}}$ (equation (5)) based on polymeric viscosity is also used to describe the influence of polymers.
@@
Given the potential for jet drops to transport pathogens or pollutants into the atmosphere, strategies to prevent their generation are pertinent. Recent studies unsurprisingly show that non-Newtonian effects, particularly that viscoplasticity and   viscoelasticity, can suppress jet drop production (Sanjay et al., 2021; Sen et al., 2021; Rodríguez-Díaz et al., 2023; Ji et al., 2023). While computational studies have successfully reproduced experimental observations, such as elasticity-induced droplet suppression (Cabalgante-Corralesa et al., 2024; Balasubramanian et al., 2024), the full impact of these effects on bubble-bursting dynamics remains elusive. In this paper, we answer the question: How does the viscoelasticity influence the observed regimes? What underlying physics governs the transitions between these regimes?   Advancements in solving non-linear constitutive equations for highly deformed interfacial flows of viscoelastic fluids   have been made possible by techniques like the log-conformation method (Fattal & Kupferman, 2004) and the square-root conformation method (Balci et al., 2011). Originally developed for single-phase flows, these methods have been extended to multiphase flows (Fraggedakis et al., 2016; López-Herrera et al., 2019; Varchanis & Tsamopoulos, 2022; França et al., 2024; Zinelis et al., 2024), facilitating more comprehensive investigations into this topic.
@@
$Oh_{p}=\frac{\eta_{p}}{\sqrt{\rho_{s}\gamma R_{0}}}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Ec\times De}},$ (5)
which is the product of $Ec$ and $De$. We note here that $Oh_{p}$ and $Oh_{s}$ are related by
$Oh_{p}=\dfrac{\eta_{p}}{\eta_{s}}\,Oh_{s}=c\,Oh_{s},$ (6)
where $c=\eta_{p}/\eta_{s}$ is the so-called concentration of the polymers (see e.g., Remmelgas et al. (1999); Hinch et al. (2024)).
[…]
In this study, we investigate viscoelastic effects on bubble bursting dynamics by exploring the three-dimensional phase space of $Oh_{s}$, $Ec$, and $De$, using volume of fluid-based finite volume simulations. Using the Oldroyd-B constitutive relation, we demonstrate   that the addition of polymers significantly influences the overall dynamics, which are governed by the interplay of viscous and elastic effects. For systems with a permanent memory of its initial state and subsequent deformations, i.e., when the additional polymeric stresses are sustained throughout the process time scale ($De\to\infty$), the dimensionless elastic modulus dictates the dynamics and suppression of jet and drops. In contrast, for systems with poor memory of its initial state and subsequent deformation ($De\to 0$), the dynamics resemble those encountered in Newtonian liquids with an effective viscosity deduced using the slender elastic jet equations. Despite its simplicity, we note that Oldroyd-B model has some crucial limitations. For instance, it cannot account for the shear-thinning behavior of polymer solutions and it predicts the divergence of stresses for strong extensional flows (Yamani & McKinley, 2023; Alves et al., 2021). Consequently, the Oldroyd-B model cannot accurately capture the final stages of filament thinning or the actual rupture of viscoelastic filaments, which may affect predictions of droplet detachment and fine aerosol formation. Nevertheless, we choose the Oldroyd-B model as its simplicity allows us to gain fundamental insight into the interplay between capillary, viscous, and elastic forces during bubble bursting.
Building upon the extensive literature on viscoelastic flows, we extend these concepts to the specific case of bubble bursting. Previous research has explored viscoelastic phenomena in various contexts, including flow through nozzles and contractions (Hinch, 1993; Chen, 1991; Boyko et al., 2024), stability and breakup of viscoelastic jets (Middleman, 1965; Goren & Gottlieb, 1982; Bousfield et al., 1986; Chang et al., 1999; Anna & McKinley, 2001; Pandey et al., 2021; Sen et al., 2024; Zinelis et al., 2024), coalescence and spreading of viscoelastic drops and bubbles (Bouillant et al., 2022; Dekker et al., 2022; Oratis et al., 2023), and oscillating bubbles in viscoelastic media (Oratis et al., 2024). Recent studies have also investigated elastoviscoplastic flows, incorporating viscous, elastic, and plastic aspects (Putz & Burghelea, 2009; Varchanis et al., 2019; França et al., 2024; Balasubramanian et al., 2024), further expanding our understanding of non-Newtonian liquids. We refer readers to reviews by Bogy (1979), Eggers (1997), and Yarin (1993) for comprehensive overviews of these topics. Our work applies the foundational knowledge developed in these works to elucidate how viscoelasticity alters the formation of Worthington jets and ejected droplets during bubble bursting, enhancing our understanding of this specific phenomenon.
@@ 2.1 Governing Equation
with $\boldsymbol{\mathcal{D}}=\left(\boldsymbol{\nabla u}+\left(\boldsymbol{\nabla u}\right)^{T}\right)/2$ representing the symmetric part of the velocity gradient tensor –equal to half of the rate-of-strain tensor.   The non-Newtonian contribution $\boldsymbol{\sigma_{p}}$ arises from the presence of polymers in the fluid. We emphasize that although we refer to $\boldsymbol{\sigma_{p}}$ as ‘polymeric stresses’ in the context of dilute polymer liquids, this concept extends to any deformable microstructure within the fluid that responds to flow (Saramito, 2007; Snoeijer et al., 2020; França et al., 2024; Balasubramanian et al., 2024). To characterize the deformation of these microstructures, we introduce the conformation tensor $\boldsymbol{\mathcal{A}}$, an order parameter that evolves from an initial identity state $\boldsymbol{\mathcal{A}}=\boldsymbol{\mathcal{I}}$ (figure 1a-ii). Here, we employ the Oldroyd-B model, which represents the simplest conformation tensor-based constitutive equation for viscoelastic fluids  (Oldroyd, 1950; Bird et al., 1977; Snoeijer et al., 2020; Stone et al., 2023; Boyko & Stone, 2024). This model assumes a linear relationship between elastic stresses and polymeric deformation,
@@
Additionally, the conformation tensor $\boldsymbol{\mathcal{A}}$ relaxes to its base state $\boldsymbol{\mathcal{I}}$ over time due to thermal effects. Once more, using the Oldroyd-B model, $\boldsymbol{\mathcal{A}}$ follows a linear relaxation law (i.e., the rate of change of $\boldsymbol{\mathcal{A}}$ in the Lagrangian frame is linear in $\boldsymbol{\mathcal{A}}$),
@@
[…]
is the frame-invariant upper convected Oldroyd derivative of second-rank tensor $\boldsymbol{\mathcal{A}}$, and $De=\lambda/\tau_{\gamma}$ (defined in equation (4)) is the Deborah number, representing the ratio of the polymer relaxation time $\lambda$ to the process timescale $\tau_{\gamma}$. We note that while the Oldroyd-B model is nonlinear in terms of the velocity field and its gradient, both the stress term and its relaxation law remain linear in $\boldsymbol{\mathcal{A}}$. This characteristic contrasts with models such as the Giesekus model, which involves a quadratic term $\boldsymbol{\mathcal{A}\cdot\mathcal{A}}$ (Giesekus, 1982), or the FENE models, which include a nonlinear term involving a finite-extensibility parameter $L$ (Bird et al., 1980). Therefore, the Oldroyd-B model is often referred to as “quasi-linear” (Davoodi et al., 2018; Alves et al., 2021) .
The Deborah number characterizes the polymeric liquid’s memory. It is instructive to note that in the limit of $De\to\infty$, polymeric liquids have permanent memory and the dissolved polymers undergo affine motion (see equation (11) and Snoeijer et al., 2020; Stone et al., 2023; Boyko & Stone, 2024)
@@
The Oldroyd-B model, despite its widespread use due to its simplicity, fails to capture several important physical phenomena (Snoeijer et al., 2020). It inadequately describes shear-thinning behavior in polymeric liquids (Yamani & McKinley, 2023) and erroneously predicts unbounded stress growth in strong extensional flows (McKinley & Sridhar, 2002; Eggers et al., 2020). These limitations can be addressed by incorporating finite polymer extension, for example, by increasing the effective $Ec$ as the polymer approaches full extension (Hinch & Harlen, 2021; Zinelis et al., 2024). Various extensions of the Oldroyd-B equations have been developed to account for such nonlinearity, either in equations (10) and (11) or in the solvent contribution in equation (9) (de Gennes, 1974; Tanner, 2000; McKinley & Sridhar, 2002; Alves et al., 2021). In this study, we employ the Oldroyd-B model to include the two primary effects of the polymer addition: the additional stress ($Ec$) and polymeric liquid memory ($De$) (Snoeijer et al., 2020). Our aim is to provide a comprehensive understanding of the entire $Ec$-$De$ parameter space (figure 1b). However, it is crucial to note that the Oldroyd-B model, while serving as a useful baseline, cannot accurately reproduce the finite-time breakup of viscoelastic filaments (Eggers et al., 2020) or the full complexity of interface rupture (Lohse & Villermaux, 2020). These limitations warrant caution when interpreting the final stages of jet thinning and droplet formation, particularly in scenarios involving strong polymer stretching.
@@ 2.2 Methods
We employ the open-source software Basilisk C (Popinet & collaborators, 2013–2024; Popinet, 2015) to solve the governing equations outlined in § 2.1. To solve the Oldroyd-B viscoelastic constitutive relation (equation (14)), Basilisk C uses the log-conformation method (Fattal & Kupferman, 2004) implemented by López-Herrera et al. (2019) which has been used extensively at finite $De$ (Turkoz et al., 2018, 2021). To explore the entire $Ec$-$De$ parameter space (figure 1c), we have extended the log-conformation formulation to solve equations (10) and (11). In the spirit of Basilisk C, this code is detailed open-source at Sanjay (2024).   The rest of the governing equations are solved using the one-fluid approximation (Tryggvason et al., 2011), with surface tension incorporated as singular body force at the liquid-gas interface (Brackbill et al., 1992). To account for the gas phase, in addition to the dimensionless parameters described in § 1 and § 2.1, we maintain constant density and viscosity ratios of $\rho_{r}=\rho_{g}/\rho_{s}=10^{-3}$ and $\eta_{r}=\eta_{g}/\eta_{s}=2\times 10^{-2}$, respectively. The liquid-gas interface is tracked using the volume of fluid (VoF) method, governed by the advection equation
@@
The initial bubble shape is determined by solving the Young-Laplace equations for quasi-static equilibrium (Princen, 1963; Toba, 1959; Villermaux et al., 2022; Sanjay, 2022). While the shape’s asymmetry increases with the Bond number $Bo$, we focus on the limit $Bo\to 0$, setting $Bo=0.001$ to regularize the singularity at the sphere-plane intersection. This results in a near-spherical initial cavity shape (figure 1a-i). We stress that here we assume that the bubble has resided at the liquid-gas interface for a duration far exceeding the polymeric medium’s relaxation time, ensuring that elasticity does not influence the initial configuration (Balasubramanian et al., 2024).   During the bubble cap bursting, the film cap retracts almost instantaneously (once again, we neglect the influence of elasticity) , after which the capillary waves are generated. As we are interested only in the bubbe cavity collapse, the simulations begin with an open cavity without the thin cap (figure 1a-ii), as also done similarly in recent studies (Deike et al., 2018; Gordillo & Rodríguez-Rodríguez, 2019; Sanjay et al., 2021). The computational domain spans $8R_{0}\times 8R_{0}$, discretized using quadtree grids with adaptive mesh refinement (AMR) (Popinet, 2009). Error tolerances for the VoF color function, curvature, velocity, and order parameter $\boldsymbol{\mathcal{A}}$ are set to $10^{-3}$, $10^{-6}$, $10^{-3}$, and $10^{-3}$, respectively.
In this work, following our earlier study (Sanjay et al., 2021), most simulations maintain a minimum grid size of $\Delta=R_{0}/512$, which dictates that, to get consistent results, 512 cells are required across the bubble radius while using uniform grids. We have also used an increased resolution ($\Delta=R_{0}/1024$ for high $De$ cases and $\Delta=R_{0}/2048$ near transitions) as needed. These resolutions are consistent with previous studies by Berny et al. (2020, 2021) on bubble bursting and Turkoz et al. (2018, 2021) on visco-elastic thinning with a maximum level of resolution of 14 (for $\Delta=R_{0}/2048$ and domain size $L_{0}=8R_{0}$). We have carried out extensive grid independence studies to ensure that changing the grid size does not influence the results (see appendix C) . We refer the readers to Popinet (2015); Sanjay (2022); Sanjay & Dixit (2024) for further details of the numerical method used in this work.